# Bidirectional Reward-Guided Diffusion for Real-World Image Super-Resolution

### v1

@@ Abstract
Diffusion-based super-resolution can synthesize rich details, but models trained on synthetic paired data often fail on real-world LR images due to distribution shifts. We propose Bird-SR, a bidirectional reward-guided diffusion framework that formulates super-resolution as trajectory-level preference optimization via reward feedback learning (ReFL), jointly leveraging synthetic LR–HR pairs and real-world LR images. For structural fidelity easily affected in ReFL, the model is directly optimized on synthetic pairs at early diffusion steps, which also facilitates structure preservation for real-world inputs under smaller distribution gap in structure levels. For perceptual enhancement, quality-guided rewards are applied at later sampling steps to both synthetic and real LR images. To mitigate reward hacking, the rewards for synthetic results are formulated in a relative advantage space bounded by their clean counterparts, while real-world optimization is regularized via a semantic alignment constraint. Furthermore, to balance structural and perceptual learning, we adopt a dynamic fidelity–perception weighting strategy that emphasizes structure preservation at early stages and progressively shifts focus toward perceptual optimization at later diffusion steps. Extensive experiments on real-world SR benchmarks demonstrate that Bird-SR consistently outperforms state-of-the-art methods in perceptual quality while preserving structural consistency, validating its effectiveness for real-world super-resolution.
Figure 1: Motivation. (a) Due to the degradation gap between synthetically generated low-resolution and real-world LR data, models trained on synthetic datasets tend to produce blurred details when applied to real-world LR images, resulting in an input distribution shift issue. (b) The TSNE visualization of Local Binary Pattern (LBP) texture features shows clearly separated clusters, indicating a significant texture distribution gap and input distribution shift between synthetic and real-world LR data. (c) Kernel density estimation (KDE) of cosine similarity distributions between low-resolution and high-resolution (HR) image pairs in high frequency band. While real-world LR images exhibit a sharp, high-similarity alignment with the ground truth (GT) characterized by a distinct multimodal distribution, synthetic LR images generated by conventional degradation models show a significant distribution broadening and a lower similarity mean. This discrepancy highlights the severe domain gap and input distribution shift that hinders the generalization of models trained solely on synthetic data.
Machine Learning, ICML
@@ 1 Introduction
In recent years, Diffusion Models (Ho et al., 2020; Song et al., 2021) have achieved remarkable progress in image generation by learning complex data distributions through progressive denoising. Building on the success of large-scale text-to-image diffusion models (Rombach et al., 2022; Esser et al., 2024; Podell et al., 2024; Labs, 2024), their strong generative priors have been increasingly adopted in image restoration tasks. In diffusion based super-resolution, the LR image serves as a conditioning signal and the reconstruction is formulated as conditional diffusion sampling, enabling pretrained diffusion models to synthesize more natural, diverse, and detailed textures than conventional supervised methods. Despite these advances, existing approaches rely heavily on synthetic paired datasets, and when applied to real world LR images often suffer from distribution shift, leading to the loss of high-frequency fine details (as visualized in Figure 1).
@@
Current diffusion based SR and reward optimization frameworks face three critical obstacles: (1) Distribution Shift Induced by Synthetic Paired Data: Training is predominantly conducted on synthetically generated LR-HR pairs, whereas real world LR degradations are far more diverse and difficult to model in Figure 1. This mismatch leads to substantial input distribution shifts, causing models to hallucinate over-smoothed or unrealistic textures. (2) Reward Hacking under Perceptual Optimization: In real-world images, the lack of paired supervision makes the model particularly prone to over-optimizing the reward signal, leading to visually implausible artifacts. (3) Trajectory-Level Trade-off in Diffusion: Finally , since different stages of the diffusion trajectory contribute differently to structure and appearance, blindly enforcing reward optimization across all steps may amplify perceptual gains at the cost of structural fidelity.
To address these challenges, we propose a bidirectional, reward-guided diffusion super-resolution framework that jointly optimizes the forward and reverse processes over both synthetic and real-world data, enabling effective alignment of real-world super-resolution capability, as illustrated in Figure 2. Our method is built upon three core ideas: (1) Directly optimization via noise injection and relative reward for synthetic data. We inject noise into clean images and directly recover them with a closed-form single-step forward formula. Leveraging the available ground truth, we enable precise supervision of structural distortion and optimize the model using relative rewards. This significantly improves optimization stability, enables flexible timestep control, and further mitigates reward hacking for real-world data. (2) Reward feedback learning for real-world low-resolution data with semantic alignment. Since the structural discrepancy between synthetic and real-world data is relatively minor at early time steps, we perform reward feedback along the reverse denoising trajectory, applying the reward signal at the later steps to optimize perceptual detail generation. Meanwhile, we incorporate semantic supervision to constrain structural consistency. (3) Trajectory-aware distortion–perception weighting along the forward diffusion trajectory. We introduce dynamic weighting to balance structural learning at early steps and perceptual learning at later steps for synthetic data.
@@
  • For synthetic data, we introduce a stable reward optimization strategy based on noise injection and relative rewards along the forward diffusion trajectory, with incorporating a trajectory-aware dynamic fidelity–perception weighting.
@@ 2.1 Diffusion based Super-resolution
Early CNN-based (Dong et al., 2016; Kim et al., 2016; Lim et al., 2017; Zhang et al., 2018b; Dai et al., 2019) and transformer-based (Zhang et al., 2022; Chen et al., 2021, 2023b; Wang and Xie, 2024; Liang et al., 2021) image super-resolution (ISR) methods focus on optimizing pixel-wise fidelity (e.g., PSNR/SSIM) via direct LR to HR mapping, but often yield overly smooth results under complex and unknown real-world degradations. To handle real-world inputs, Real-ISR approaches simulate diverse degradation processes during training, with representative methods including BSRGAN (Zhang et al., 2021) and Real-ESRGAN (Wang et al., 2021), while GAN-based methods (Chan et al., 2021; Chen et al., 2023a; Liang et al., 2022) such as ESRGAN (Wang et al., 2018) further enhance perceptual quality through adversarial learning.
@@
Recent diffusion-based ISR methods (Cui et al., 2024a; Dong et al., 2025; Wang et al., 2024b; Chen et al., 2025a; Wu et al., 2024a; Chen et al., 2025b; Cheng et al., 2025; Wan et al., 2025; Sun et al., 2024b) leverage the strong generative priors of pre-trained text-to-image (T2I) diffusion models (Rombach et al., 2022; Esser et al., 2024; Podell et al., 2024; Labs, 2024) to address Real-ISR. StableSR (Wang et al., 2024a) injects LR structural information into Stable Diffusion via ControlNet (Zhang et al., 2023a), while DiffBIR (Lin et al., 2024) and PASD (Yang et al., 2024) adopt degradation-restoration pipelines followed by conditional diffusion-based detail enhancement. Beyond pixel-level conditioning, SeeSR (Wu et al., 2024b), CoSeR (Sun et al., 2024a), and PiSA-SR (Sun et al., 2025) introduce language–vision collaboration to enable semantic-aware and controllable diffusion. From an efficiency perspective, ResShift (Yue et al., 2023) embeds LR information into the initial noise space, where as InvSR (Yue et al., 2025) and UPSR (Zhang et al., 2025) further refine noise prediction and inversion strategies. Scaling model and data size, as demonstrated by SUPIR (Yu et al., 2024) and DreamClear (Ai et al., 2024), significantly improves photorealism, and DiT4SR (Duan et al., 2025) introduces the Diffusion Transformer (Peebles and Xie, 2023) (DiT) into SR by incorporating an LR-conditioned stream within DiT blocks.
@@
Figure 2: Overview of the Bird-SR, a bidirectional reward-guided diffusion framework for real-world Super-Resolution. For synthetic low-resolution data, predefined noise is injected into clean images $x_{0}^{1:n}$, and intermediate predictions $\hat{x}_{0}^{1:n}$ are obtained via closed-form single-step interpolation. For real-world low-resolution data, sampling starts from pure noise $x_{T}^{1:n}$ and optimizes only the final timestep along the reverse diffusion trajectory, with a reference model output $\tilde{x}_{0}^{1:n}$ introduced to enforce semantic alignment.
2.2 Preferences Optimization for Diffusion Model
@@
Recent studies have explored aligning diffusion models with human preferences or task-specific objectives through preference optimization. One line of work focuses on direct fine-tuning with scalar reward signals, where differentiable rewards are backpropagated through the denoising process. Representative methods include ReFL (Xu et al., 2023), DRaFT (Clark et al., 2023), and AlignProp (Prabhudesai et al., 2023), which inject reward feedback into diffusion training to guide perceptual or semantic improvements, while employing truncated backpropagation or memory-efficient techniques to alleviate optimization instability. Another line of research aligns diffusion models from a reinforcement learning perspective using policy gradient–based methods. DDPO (Black et al., 2024) and DPOK (Fan et al., 2023) treat the reverse diffusion process as a sequential decision trajectory and optimize model parameters via expected rewards. Diffusion-DPO (Wallace et al., 2023) extends direct preference optimization to diffusion models using pairwise preferences, avoiding explicit reward modeling. More recently, GRPO-based strategies (Shao et al., 2024) improve training stability and scalability, with Flow-GRPO (Liu et al., 2025) and Dance-GRPO (Xue et al., 2025) demonstrating effectiveness for large-scale visual generation and preference alignment. SRPO (Shen et al., 2025) introduces Direct-Align, which predefines a noise prior and formulates rewards as text-conditioned signals. In contrast to prior approaches, our method adopts a mixed forward–backward optimization framework that enables stable and timestep-aware reward feedback, leading to improved robustness in real-world image super-resolution.
@@ 3 Preliminary
3.1 Diffusion Models and Conditional Diffusion for SR
Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020; Song et al., 2021) are generative models that learn complex data distributions through a progressive denoising process. Given a high-resolution image $x_{0}$, the forward diffusion process gradually adds Gaussian noise over $T$ timesteps:
@@ 4 Method
In this section, we present Bird-SR, a reward–guided diffusion framework for real-world image super-resolution, as illustrated in Figure 2. The key idea of our approach is to align diffusion-based super-resolution models with the trajectory-level preferences, jointly leveraging synthetic LR–HR pairs and real-world LR images through stable bidirectional reward feedback.
@@ 4.1 Forward Optimization with Relative Reward
For synthetic paired data, we adopt a directly injected noise optimization strategy. For forward processes, instead of performing full diffusion sampling, we inject predefined gaussian noise into clean high-resolution images at a randomly selected timestep. Given a clean high-resolution image $x_{0}$, the low-resolution image $y$, and the forward diffusion process injects gaussian noise at timestep $t$ as Equ. 1, then initialize model with a pre-trained conditional denoising network $\epsilon_{\theta}(x_{t},t\mid y)$. Instead of performing multi-step reverse diffusion, we reconstruct $x_{0}$ using a closed-form single-step interpolation:
@@
where $r$ denotes a reward function that measures the perceptual quality of an image, and $\phi$ denotes preference loss function. In addition, perceptual learning on high-resolution images facilitates optimization for real-world data.
@@ 4.2 Reward Feedback with Semantic Alignment
where the estimate $\hat{x}_{0}$ is obtained via Eq. (3), $r$ denotes a reward function that measures the perceptual quality of an image, and $\phi$ denotes preference loss function. We focus on the last timestep for reward supervision because early timesteps mainly define global structure, which is already well-aligned with real-world data, and only the prediction at last timestep is optimized.
While reward-guided optimization encourages perceptually favorable outputs, only applying rewards optimization often leads to reward hacking, especially for real-world LR images with complex and unknown degradation patterns. To address this issue, we incorporate semantic alignment into the reward feedback learning process. Specifically, for real-world LR images, we leverage DINO-based (Oquab et al., 2023) spatial semantic features to guide the reward guided reverse diffusion process:
@@
4:  Model Predict $\hat{x}_{0}=(x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},t\mid y^{\mathrm{syn}}))/{\alpha_{t}}$
@@
10:  Sample $(x,y)\sim\mathcal{D}_{\mathrm{real}}$
11:  Reverse from $t=T$ to $t=1$ for $\hat{x}_{0}$
12:  Last timestep state $x_{1}=\alpha_{1}\text{stop\_grad}(\hat{x}_{0})+\sigma_{1}\epsilon$
13:  Last predict $\hat{x}_{0}=(x_{1}-\sigma_{1}\epsilon_{\theta}(x_{1},1\mid y^{\mathrm{real}}))/\alpha_{1}$
14:  Reference predict from $\epsilon_{\text{ref}}\longrightarrow{\tilde{x}_{0}}$
15:  Absolute reward $r(\hat{x}_{0})$
16:  $\mathcal{L}_{\text{unpair}}=\mathbb{E}_{t}\big[\phi(r(\hat{x}_{0}))\big]$ (Eq. 7)
@@
18:  $\mathcal{L}_{\text{reverse}}=\mathcal{L}_{\text{unpair}}(\theta)+\lambda_{\text{sem}}\mathcal{L}_{\text{sem-align}}(\theta)$ (Eq. 9)
19:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{reverse}}$ {Gradient update reverse}
20:  end for
21:  return $\theta$
Figure 3: Qualitative comparisons with state-of-the-art Real-ISR methods. Our method performs best in terms of image realism and detail generation especially preserving fine structures and restoring text details. More visual results can be found in the Appendix.
4.3 Dynamic Distortion-perception Weight

### v2

# Bird-SR: Bidirectional Reward-Guided Diffusion for Real-World Image Super-Resolution

@@ Abstract
Powered by multimodal text-to-image priors, diffusion-based super-resolution excels at synthesizing intricate details; however, models trained on synthetic low-resolution (LR) and high-resolution (HR) image pairs often degrade when applied to real-world LR images due to significant distribution shifts. We propose Bird-SR, a bidirectional reward-guided diffusion framework that formulates super-resolution as trajectory-level preference optimization via reward feedback learning (ReFL), jointly leveraging synthetic LR-HR pairs and real-world LR images. For structural fidelity easily affected in ReFL, the model is directly optimized on synthetic pairs at early diffusion steps, which also facilitates structure preservation for real-world inputs under smaller distribution gap in structure levels. For perceptual enhancement, quality-guided rewards are applied to both synthetic and real LR images at the later trajectory phase. To mitigate reward hacking, the rewards for synthetic results are formulated in a relative advantage space bounded by their ground-truth counterparts, while real-world optimization is regularized via a semantic alignment constraint. Furthermore, to balance structural and perceptual learning, we introduce a dynamic fidelity-perception weighting strategy that emphasizes structure preservation at early stages and progressively shifts focus toward perceptual optimization at later diffusion steps. Extensive experiments on real-world SR benchmarks demonstrate that Bird-SR consistently outperforms state-of-the-art methods in perceptual quality while preserving structural consistency, validating its effectiveness for real-world super-resolution. Our code can be obtained at https://github.com/fanzh03/Bird-SR.
@@ 1 Introduction
Figure 1: Motivation. (a) Due to the degradation gap between synthetically generated low-resolution and real-world low-resolution data, models trained on synthetic datasets tend to produce blurred details when applied to real-world low-resolution images, resulting in an input distribution shift issue. (b) The t-SNE visualization of Local Binary Pattern (LBP) texture features shows clearly separated clusters, indicating a significant texture distribution gap and input distribution shift between synthetic and real-world LR data. (c) Kernel density estimation (KDE) of cosine similarity distributions between low-resolution and high-resolution image pairs in high frequency band. While real-world LR images exhibit a sharp, high-similarity alignment with the ground truth (GT) characterized by a distinct multimodal distribution, synthetic LR images generated by conventional degradation models show a significant distribution broadening and a lower similarity mean. This discrepancy highlights the severe domain gap and input distribution shift that hinders the generalization of models trained solely on synthetic data.
In recent years, Diffusion Models [DDPM, song2021scorebased] have achieved remarkable progress in image generation by learning complex data distributions through progressive denoising. Building on the success of large-scale text-to-image diffusion models [sd, sd3, sdxl, flux, fang2024vividvideovirtualtryon], their strong generative priors have been increasingly adopted in image restoration tasks. In diffusion based super-resolution, the LR image serves as a conditioning signal and the reconstruction is formulated as conditional diffusion sampling, enabling pretrained diffusion models to synthesize more natural, diverse, and detailed textures than conventional supervised methods. Despite these advances, existing approaches rely heavily on synthetic paired datasets, and when applied to real world LR images, they often suffer from distribution shift, leading to the loss of high-frequency fine details (as visualized in Figure 1).
@@
Figure 2: Evolution of semantic and texture feature spaces during the reverse diffusion process. We visualize the t-SNE of intermediate predictions ($\hat{x}_{0}$) from real (red) and synthetic (cyan) reverse trajectories across early, middle, and late denoising stages. Top: VGG features demonstrate that macroscopic semantic structures remain highly consistent throughout the entire process. Bottom: LBP features reveal that, compared to semantic features, texture information presents a distinct domain gap. Moreover, as the reverse timestep progresses, the domains become increasingly dispersed.
Current diffusion based SR and reward optimization frameworks face three critical obstacles: (1) Distribution Shift Induced by Synthetic Paired Data: Training is predominantly conducted on synthetically generated LR-HR pairs, whereas real world LR degradations are far more diverse and difficult to model, as illustrated in Figure 1. This mismatch leads to substantial input distribution shifts, causing models to hallucinate over-smoothed or unrealistic textures. (2) Reward Hacking under Perceptual Optimization: In real-world images, the lack of paired supervision makes the model particularly prone to over-optimizing the reward signal, leading to visually implausible artifacts. (3) Trajectory-Level Trade-off in Diffusion: Finally, since different stages of the diffusion trajectory contribute differently to structure and appearance, blindly enforcing reward optimization across all steps may amplify perceptual gains at the cost of structural fidelity.
To address these challenges, we propose a bidirectional reward-guided diffusion super-resolution framework that jointly optimizes the forward and reverse processes over both synthetic and real-world data, enabling effective alignment of real-world SR capability, as illustrated in Figure 3. Our method is built upon three core ideas: (1) Direct optimization via noise injection and relative reward for synthetic data. We inject noise into clean images and directly recover them with a closed-form single-step forward formula. Leveraging the available ground truth, we enable precise supervision of structural distortion and optimize the model using relative rewards. This significantly improves optimization stability, enables flexible timestep control, and further mitigates reward hacking for real-world data. (2) Reward feedback learning for real-world low-resolution data with semantic alignment. Figure 2 shows that during the reverse denoising process, while the semantic representations of synthetic and real data are highly aligned, their high-frequency perceptual details manifest a pronounced domain divergence. Motivated by this observation, we perform reward feedback along the reverse denoising trajectory, applying the reward signal at the later steps to optimize perceptual detail generation. This naturally corresponds to the temporal mechanism of diffusion models, which generate semantics in the early stages and synthesize textures in the later stages. Meanwhile, we incorporate semantic supervision to constrain structural consistency. (3) Trajectory-aware distortion-perception weighting along the forward diffusion trajectory. We introduce dynamic weighting to balance structural learning at early steps and perceptual learning at later steps for synthetic data.
@@
  • For synthetic data, we introduce a stable reward optimization strategy based on noise injection and relative rewards along the forward diffusion trajectory, by incorporating a trajectory-aware dynamic fidelity-perception weighting.
@@ 2.1 Diffusion based Super-Resolution
Early CNN-based [SRCNN, VDSR, EDSR, RCAN, 8954252] and transformer-based [10.1007/978-3-031-19790-1_39, IPT, chen2023activating, 10824429, 9607618] image super-resolution (ISR) methods focus on optimizing pixel-wise fidelity (e.g., PSNR/SSIM) via direct LR to HR mapping, but often yield overly smooth results under complex and unknown real-world degradations. To handle real-world inputs, Real-ISR approaches simulate diverse degradation processes during training, with representative methods including BSRGAN [BSRGAN] and Real-ESRGAN [realesrgan], while GAN-based methods [GLEAN, HGGT, 10.1007/978-3-031-19797-0_33] such as ESRGAN [esrgan] further enhance the perceptual quality through adversarial learning.
@@
Recent diffusion-based ISR methods [DoSSR, TSD-SR, SinSR, AdcSR, OSEDiff, FaithDiff, cheng2025effective, ControlSR, sun2023ccsr] leverage the strong generative priors of pre-trained text-to-image (T2I) diffusion models [sd, sd3, sdxl, flux] to address Real-ISR. StableSR [StableSR] injects LR structural information into Stable Diffusion via ControlNet [ControlNet], while DiffBIR [DiffBIR] and PASD [PASD] adopt degradation-restoration pipelines followed by conditional diffusion-based detail enhancement. Beyond pixel-level conditioning, SeeSR [SeeSR], CoSeR [CoSeR], and PiSA-SR [PiSA-SR] introduce language and vision collaboration to enable semantic-aware and controllable diffusion. From an efficiency perspective, ResShift [ResShift] embeds LR information into the initial noise space, whereas InvSR [InvSR] and UPSR [UPSR] further refine noise prediction and inversion strategies. Scaling model and data size, as demonstrated by SUPIR [SUPIR] and DreamClear [DreamClear], significantly improves photorealism. DiT4SR [DiT4SR] introduces the Diffusion Transformer [dit] (DiT) into SR by incorporating an LR-conditioned stream within DiT blocks.
@@
Figure 3: Overview of the proposed Bird-SR, a bidirectional reward-guided diffusion framework for real-world super-resolution. For synthetic low-resolution data, predefined noise is injected into clean images $x_{0}^{1:n}$, and intermediate predictions $\hat{x}_{0}^{1:n}$ are obtained via closed-form single-step interpolation. For real-world low-resolution data, sampling starts from pure noise $x_{T}^{1:n}$ and optimizes only the final timestep along the reverse diffusion trajectory, with a reference model output $\tilde{x}_{0}^{1:n}$ introduced to enforce semantic alignment.
2.2 Preferences Optimization for Diffusion Models
@@
Recent studies have explored aligning diffusion models with human preferences or task-specific objectives through preference optimization. One line of work focuses on direct fine-tuning with scalar reward signals, where differentiable rewards are backpropagated through the denoising process. Representative methods include ReFL [refl], DRaFT [draft], and AlignProp [alignprop], which inject reward feedback into diffusion training to guide perceptual or semantic improvements, while employing truncated backpropagation or memory-efficient techniques to alleviate optimization instability. Another line of research aligns diffusion models from a reinforcement learning perspective using policy gradient–based methods. DDPO [ddpo] and DPOK [dpok] treat the reverse diffusion process as a sequential decision trajectory and optimize model parameters via expected rewards. Diffusion-DPO [diffusion-dpo] extends direct preference optimization to diffusion models using pairwise preferences, avoiding explicit reward modeling. More recently, GRPO-based strategies [grpo] improve training stability and scalability, with Flow-GRPO [flow-grpo] and Dance-GRPO [dancegrpo] demonstrating effectiveness for large-scale visual generation and preference alignment. SRPO [srpo] introduces Direct-Align, which predefines a noise prior and formulates rewards as text-conditioned signals. In contrast to prior approaches, our method adopts a mixed forward–backward optimization framework that enables stable and timestep-aware reward feedback, leading to improved robustness in real-world image super-resolution.
@@ 3 Preliminary
3.1 Conditional Diffusion for Super-Resolution
Denoising Diffusion Probabilistic Models (DDPMs) [DDPM, song2021scorebased] are generative models that learn complex data distributions through a progressive denoising process. Given a high-resolution image $x_{0}$, the forward diffusion process gradually adds gaussian noise over $T$ timesteps:
@@ 4 Method
In this section, we present Bird-SR, a reward-guided diffusion framework for real-world image super-resolution, as illustrated in Figure 3. The key idea of our approach is to align diffusion-based super-resolution models with the trajectory-level preferences, jointly leveraging synthetic LR-HR pairs and real-world LR images through stable bidirectional reward feedback.
@@ 4.1 Forward Optimization with Relative Reward
For synthetic paired data, we adopt a directly injected noise optimization strategy. For forward processes, instead of performing full diffusion sampling, we inject predefined gaussian noise into clean high-resolution images at a randomly selected timestep. Given a clean high-resolution image $x_{0}$ and its corresponding low-resolution image $y$, with the forward diffusion process injecting gaussian noise at timestep $t$ as defined in Eq. 1, the model is initialized using a pre-trained conditional denoising network $\epsilon_{\theta}(x_{t},t\mid y)$. Instead of performing multi-step reverse diffusion, we reconstruct $x_{0}$ using a closed-form single-step interpolation:
@@
where $r$ denotes a reward function that measures the perceptual quality of an image, and $\phi$ denotes the preference loss function. In addition, perceptual learning on high-resolution images facilitates optimization for real-world data.
@@ 4.2 Reward Feedback with Semantic Alignment
where the estimate $\hat{x}_{0}$ is obtained via Eq. (3), $r$ denotes a reward function that measures the perceptual quality of an image, and $\phi$ denotes preference loss function. We focus on the last timestep for reward supervision because early timesteps mainly define global structure, which is already well-aligned with real-world data as shown in Figure 2, and only the prediction at last timestep is optimized.
While reward-guided optimization encourages perceptually favorable outputs, only applying reward optimization often leads to reward hacking, especially for real-world LR images with complex and unknown degradation patterns. To address this issue, we incorporate semantic alignment into the reward feedback learning process. Specifically, for real-world LR images, we leverage DINO-based [oquab2023dinov2] spatial semantic features to guide the reward guided reverse diffusion process:
@@
Figure 4: Qualitative comparisons with state-of-the-art Real-ISR methods. Our method performs best in terms of image realism and detail generation especially preserving fine structures and restoring text details. More visual results are in the Appendix.
4.3 Dynamic Distortion-perception Weight
@@
4:   Model Predict $\hat{x}_{0}=(x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},t\mid y^{\mathrm{syn}}))/{\alpha_{t}}$ (Eq. 5)
@@
10:   Sample $y\sim\mathcal{D}_{\mathrm{real}}$, $x\sim\mathcal{N}(0,I)$
11:   for $t=T$ to $t=j+2$ do
12:    no grad: reverse for $x_{t-1}\leftarrow x_{t}$
13:   end for
14:   with grad: $x_{j}\leftarrow x_{j+1}$ (Eq. 3)
15:   Last predict $\hat{x}_{0}\leftarrow x_{j}$
16:   Reference predict from $\epsilon_{\text{ref}}\longrightarrow{\tilde{x}_{0}}$
17:   Absolute reward $r(\hat{x}_{0})$
18:   $\mathcal{L}_{\text{unpair}}=\mathbb{E}_{t}\big[\phi(r(\hat{x}_{0}))\big]$ (Eq. 7)
@@
20:   $\mathcal{L}_{\text{reverse}}=\mathcal{L}_{\text{unpair}}(\theta)+\lambda_{\text{sem}}\mathcal{L}_{\text{sem-align}}(\theta)$ (Eq. 9)
21:   Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{reverse}}$ {Gradient update reverse}
22:  end for
23:  return $\theta$